# ReDeLEx: A Framework for Relational Deep Learning Exploration

### v1

@@ Abstract
Keywords:
Relational Deep Learning Relational Databases Graph Neural Networks
@@ 1 Introduction
From their establishment [9], Relational Databases (RDBs) played a pivotal role in transforming our society into the current information age. Data stored as interconnected tables, safeguarded by integrity constraints, have proven to be an effective method for managing domain information. Consequently, RDBs still prevail today as a backbone of critical systems in a number of important domains ranging from healthcare [41] to government [28].
@@
Although ubiquitous in modern application stacks, the data format of RDBs is deeply incompatible with classic Machine Learning (ML) workflows, which assume data in the standard form of fixed-size i.i.d. feature vectors, forming the common “tabular” learning format. Nevertheless, this assumption is clearly violated with the relationships between the differently-sized RDB tables. To address the discrepancy, the historically prevailing approach has been to turn the relational into the tabular format by means of “propositionalization” [25], which is essentially a feature extraction routine where relational substructures get aggregated from the relations into the attributes (features) of the tabular format, upon which classical ML methods may then operate. Nevertheless, this comes at the cost of information loss during this preprocessing step.
@@
Recently, building on advances in graph representation learning [19], deep learning models directly exploiting the relational structure of RDBs have started to gain traction [13, 44, 43, 32], establishing the field of Relational Deep Learning (RDL) [15]. Following the “message-passing” principles of Graph Neural Networks (GNN; [42]), RDL models treat the structure of an RDB as a heterogeneous (temporal) graph, where individual table rows correspond to nodes, and edges are formed through integrity constraints set by the primary and foreign keys. Utilizing the graph representation then allows for the application of various GNNs, and their various extensions, with adapted message-passing schemes.
@@
The generality and spread of RDBs allow for a broad spectrum of domain information to be stored, upon which a variety of predictive tasks can be formulated, each with unique aspects and qualities. This presents a challenge for establishing a broad enough benchmark to appropriately assess the general performance of RDL. Currently, the most prominent effort in this area is the recently proposed RelBench [33], which introduced the evaluation of RDL, albeit with a very limited scope of simple models and just five accessible datasets. However, the overarching domain of relational learning [14], currently ignored by the RDL community, has a rich history of working with the relational data format [30, 12], including benchmarking of the propositionalization techniques [25]. Notably, this includes the CTU Relational Learning Repository [29] that historically collected more than 70 diverse RDBs.
@@
Our aim in this paper is to provide a bridge between the communities of traditional (logic-based) relational learning [14] and the contemporary RDL [33] towards a more comprehensive evaluation of the diverse existing methods. To that aim, we introduce ReDeLEx—an experimental framework for developing and benchmarking diverse RDL architectures against classic methods over the most comprehensive collection of tasks and datasets to date. The implementation of the framework is readily available on GitHub. [^1] https://github.com/jakubpeleska/ReDeLEx
@@ 2.1 Relational Databases
Principles of RDBs are formally based on the relational model [10], which is grounded in relational logic [16]. This abstraction enables the definition of any database, regardless of specific software implementation, as a collection of $n$-ary relations, which are defined over the domains of their respective attributes, managed by the Relational Database Management System (RDBMS) to ensure data consistency with the integrity constraints of the database schema. The key concepts to be used in this paper are as follows.
@@ 2.1.1 Relational Database
A Relational Database (RDB) $\mathcal{R}$ is defined as a finite set of relations $R_{1},R_{2},\dots,R_{n}$. An instance of an RDB $\mathcal{R}$ is implemented through a RDBMS, enabling to perform Structured Query Language (SQL; [5]) operations, rooted in relational algebra.
@@ 2.1.5 Integrity constraints
In addition to the domain constraints $\mathsf{dom}(A_{i})$, the most important integrity constraints are the primary and foreign keys. A primary key $PK$ of a relation $R$ is a minimal subset of its attributes $R[PK]\subseteq\mathcal{A_{R}}$ that uniquely identifies each tuple: $\forall t_{1},t_{2}\in R:~{}(t_{1}[PK]=t_{2}[PK])\Rightarrow(t_{1}=t_{2})$. A foreign key ${FK}_{R_{2}}$ in relation $R_{1}$ then refers to the primary key ${PK}$ of another relation $R_{2}$ as $\forall t\in R_{1}:~{}t[FK]\in\{t^{\prime}[PK]\mid t^{\prime}\in R_{2}\}\,.$ This constitutes the inter-relations in the database, with the RDBMS handling the referential integrity of ${T_{R_{1}}}[FK]\subseteq{T_{R_{2}}}[PK]$.
@@ 2.2 Graph Neural Networks
Graph Neural Networks constitute a comprehensive class of neural models designed to process graph-structured data through the concept of (differentiable) message-passing [42]. Given an input graph $G=(\mathcal{V},\mathcal{E})$, with a set of nodes $\mathcal{V}$ and edges $\mathcal{E}$, let $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}}$ be the vector representation (embedding) of node $v$ at layer $l$. The general concept of GNNs can then be defined through the following sequence of three functions:
@@ 2.3 Relational Deep Learning
In this paper, we adopt the concept of RDL as extending mainstream deep learning models, particularly the GNNs (Sec. 2.2), for application to RDBs (Sec. 2.1). For completeness, in the relational learning community [12], a number of similar approaches combining relational (logic-based) and deep learning methods arose under a similar name of “deep relational learning” [35]. Nevertheless, for compatibility with the recently introduced frameworks [15], we hereby continue with the contemporary RDL view, where RDBs are first transformed into a graph-based representation suitable for the GNN-based learning.
@@ 2.3.1 Database Representation
The fundamental characteristic of RDL [15] is to represent an RDB as a heterogeneous graph.222sometimes referred to as the “relational entity graph” The graph representation can be defined as $G=(\mathcal{V},\mathcal{E},\mathcal{T}^{v},\mathcal{T}^{e})$, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}$ is the set of edges, $\mathcal{T}^{v}$ is a set of node types with a mapping $\phi:\mathcal{V}\to\mathcal{T}^{v}$, and $\mathcal{T}^{e}$ is a set of edge types with a mapping $\psi:\mathcal{E}\to\mathcal{T}^{e}$. The node types and edge types collectively form the graph schema $(\mathcal{T}^{v},\mathcal{T}^{e})$.
Given an RDB schema $\mathcal{R}$, the node types $T\in\mathcal{T}^{v}$ correspond to the relations (tables) $T$ within the database $\mathcal{T}^{v}\overset{1:1}{\to}\mathcal{R}$, while the edge types $\mathcal{T}^{e}$ represent the undirected inter-relations between the tables, as defined by the primary-foreign key pairs: $\mathcal{T}^{e}=\{({R_{i},R_{j}})~{}|~{}{R_{i}}[FK_{R_{j}}]\subseteq{R_{j}}[PK]~{}\lor~{}{R_{j}}[FK_{R_{i}}]\subseteq{R_{i}}[PK]\}\text{.}$ For a specific instance of an RDB $\mathcal{R}$, the set of nodes $\mathcal{V}$ is then defined as the union of all tuples (rows) $t_{i}$ from each relation $\mathcal{V}=\{v_{i,j}~{}|~{}R_{i}\in\mathcal{R},~{}t_{j}\in R_{i}\}\text{,}$ and the set of edges $\mathcal{E}$ is defined as =$\mathcal{E}=\{({v_{i,k},v_{j,l}})|~{}t_{k}\in R_{i},~{}t_{l}\in R_{j},(R_{i},R_{j})\in\mathcal{T}^{e}\}$.
The graph representation is further enriched by node embedding matrices, attribute schema, and optionally a time mapping. Node embedding matrix $h^{(l)}_{v}\in\mathbb{R}^{d\times d_{\phi(v)}}$ contains the embedding representation of a node $v\in\mathcal{V}$ in a given layer $l$. With an attribute schema $\mathcal{A}_{T}$ that provides information about the types of attributes $A_{1},\dots,A_{n}$ associated with the nodes $v$ of a specific node type $T\in\mathcal{T}^{v}$, the initial embedding tensors $h_{v}^{(0)}\in\mathbb{R}^{d^{(0)}\times n}$ are computed from the raw database attribute tuples ${t_{i}}=(a_{1},a_{2},\ldots,a_{n})$ through multi-modal attribute encoders [15]. Finally, the time mapping is a function $\tau$ that assigns a timestamp $t_{v}$ to each node $\tau:v\mapsto t_{v}$, effectively creating a dynamically growing graph in time, enabling the use of temporal graph sampling [34].
@@ 2.3.2 Predictive Tasks
In RDL, predictive tasks are implemented through the creation of dedicated training tables $T_{t}$ that extend the existing relational schema of $\mathcal{R}$. As introduced in [15], a training table $T_{t}$ contains two essential components: foreign keys $T_{t}[FK]$ that identify the entities of interest and target labels $y\in\mathcal{A}_{T_{t}}\setminus T_{t}[FK]$. Additionally, timestamps $t_{v}\in\mathcal{A}_{T_{t}}$ that define temporal boundaries for the prediction of $y$ can also be included.
@@ 2.3.3 Neural Architecture Space
    Table-level tabular model allows to employ existing tabular learning models [6, 21] to yield more sophisticated node embeddings $h_{v}^{(l)}$. Notably, in this stage, an RDL model may reduce the dimensionality of the node attribute matrix embedding $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}\times n}$ to a vector embedding $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}_{\phi(v)}}$.
@@
    Graph neural model then depends on the chosen embedding dimensionality of $h_{v}^{(l)}$. If there is a single embedding vector $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}_{\phi(v)}}$ per each node, the model can employ standard GNN (Sec. 2.2) heterogeneous message-passing [38, 4], otherwise a custom message-passing scheme [32] is required.
@@ 3.1 Workflow Components
The ReDeLEx workflow, depicted in Fig. 1, consists of modular blocks that enable systematic exploration of the neural architecture and database configuration space, significantly extending the current scope [33] of RDL experimentation.
@@ 3.1.1 Database Connectivity
In contrast to [33], the framework provides a standardized interface for connecting directly to an RDB [2], supporting various dialects of RDBMS. Notably, this enables a truly end-to-end deep learning pipeline, connecting to a possibly remote RDBMS, hosting the target RDB.
@@ 3.1.3 Predictive Tasks
The existing benchmark [33] provides support solely for tasks with a training table $T_{t}$ generated from historical data through an SQL query. While useful, without any changes to the underlying database, this setting renders many RDB prediction tasks infeasible. ReDeLEx addresses this problem by adding support for tasks that require more substantial modifications of the original database. Tasks leveraging this functionality then not only generate a new table $T_{t}$ but a whole modified instance $\mathcal{R}^{\prime}$ of the original database $\mathcal{R}$.
@@
Importantly, this scheme can be applied to generate tasks for unsupervised pretraining. Pretraining tasks can be created by choosing any table $T\in\mathcal{R}$ in the database and duplicating it as $T^{\prime}$. The unchanged duplicate $T^{\prime}$ can then be used as a training table $T_{t}$, while the values of cells in the original table $T$ are randomly removed (masked out). The task is then to reconstruct any missing values in the classical tabular learning fashion [1], opening possibilities for sophisticated pretraining methods [36].
@@ 3.2 RDL-Suitable Databases and Tasks
Due to the generality of the relational model, RDBs often contain data with vastly diverse structural characteristics that, in some cases, do not properly exploit the relational model (Sec. 2.1). Likewise, not all of the 70+ available RDBs [29] are actually suitable for the relational learning models. In this section, we examine RDB characteristics in the context of RDL to identify suitable databases to be used in the experiments (Sec. 4).
@@ 3.2.3 Graph Data
On the other hand, RDBs are also characterized by building on the tabular representation, where an arbitrary number of attributes can be connected by a single relation. This is in contrast to the graph data which correspond to binary relational structures. Consequently, natively graph-structured data, such as molecules or family trees, although possible to be stored in an RDB, also do not fully exploit the relational model. In such cases, the RDL paradigm reduces to the simpler GNN setting [19], introducing an unnecessary complexity otherwise. More generally, RDL models for tasks on RDBs with a low number of non-key attributes (see Sec. 4.3) may suffer from information sparsity (see App. Table 5 for databases with the stated characteristics.)

### v2

@@ 1 Introduction
From their establishment [Codd1970], Relational Databases (RDBs) played a pivotal role in transforming our society into the current information age. Data stored as interconnected tables, safeguarded by integrity constraints, have proven to be an effective method for managing domain information. Consequently, RDBs still prevail today as a backbone of critical systems in a number of important domains ranging from healthcare [white_pubmed_2020] to government [maali_enabling_2010].
@@
Although ubiquitous in modern application stacks, the data format of RDBs is deeply incompatible with classic Machine Learning (ML) workflows, which assume data in the standard form of fixed-size i.i.d. feature vectors, forming the common “tabular” learning format. Nevertheless, this assumption is clearly violated with the relationships between the differently-sized RDB tables. To address the discrepancy, the historically prevailing approach has been to turn the relational into the tabular format by means of “propositionalization” [propos], which is essentially a feature extraction routine where relational substructures get aggregated from the relations into the attributes (features) of the tabular format, upon which classical ML methods may then operate. Nevertheless, this comes at the cost of information loss during this preprocessing step.
@@
Recently, building on advances in graph representation learning [hamilton_graph_2020], deep learning models directly exploiting the relational structure of RDBs have started to gain traction [Cvitkovic2020, zhang2023gfs, zahradnik2023deep, peleska_transformers_2024], establishing the field of Relational Deep Learning (RDL) [fey2024position]. Following the “message-passing” principles of Graph Neural Networks (GNN; [wu2020comprehensive]), RDL models treat the structure of an RDB as a heterogeneous (temporal) graph, where individual table rows correspond to nodes, and edges are formed through integrity constraints set by the primary and foreign keys. Utilizing the graph representation then allows for the application of various GNNs, and their various extensions, with adapted message-passing schemes.
@@
The generality and spread of RDBs allow for a broad spectrum of domain information to be stored, upon which a variety of predictive tasks can be formulated, each with unique aspects and qualities. This presents a challenge for establishing a broad enough benchmark to appropriately assess the general performance of RDL. Currently, the most prominent effort in this area is the recently proposed RelBench [robinson2024relbench], which introduced the evaluation of RDL, albeit with a very limited scope of simple models and just five accessible datasets. However, the overarching domain of relational learning [Raedt], currently ignored by the RDL community, has a rich history of working with the relational data format [muggleton1994inductive, cropper2020turning30], including benchmarking of the propositionalization techniques [propos]. Notably, this includes the CTU Relational Learning Repository [motl2015ctu] that historically collected more than 70 diverse RDBs.
@@
Our aim in this paper is to provide a bridge between the communities of traditional (logic-based) relational learning [Raedt] and the contemporary RDL [robinson2024relbench] towards a more comprehensive evaluation of the diverse existing methods. To that aim, we introduce ReDeLEx—an experimental framework for developing and benchmarking diverse RDL architectures against classic methods over the most comprehensive collection of tasks and datasets to date. The implementation of the framework is readily available on GitHub. [^1] https://github.com/jakubpeleska/ReDeLEx
@@ 2.1 Relational Databases
Principles of RDBs are formally based on the relational model [codd1990relational], which is grounded in relational logic [gallier2015logic]. This abstraction enables the definition of any database, regardless of specific software implementation, as a collection of $n$-ary relations, which are defined over the domains of their respective attributes, managed by the Relational Database Management System (RDBMS) to ensure data consistency with the integrity constraints of the database schema. The key concepts to be used in this paper are as follows.
@@ 2.1.1 Relational Database
A Relational Database (RDB) $\mathcal{R}$ is defined as a finite set of relations $R_{1},R_{2},\dots,R_{n}$. An instance of an RDB $\mathcal{R}$ is implemented through a RDBMS, enabling to perform Structured Query Language (SQL; [chamberlin_sequel_1974]) operations, rooted in relational algebra.
@@ 2.1.5 Integrity constraints
In addition to the domain constraints $\mathsf{dom}(A_{i})$, the most important integrity constraints are the primary and foreign keys. A primary key $PK$ of a relation $R$ is a minimal subset of its attributes $R[PK]\subseteq\mathcal{A_{R}}$ that uniquely identifies each tuple: $\forall t_{1},t_{2}\in R:~(t_{1}[PK]=t_{2}[PK])\Rightarrow(t_{1}=t_{2})$. A foreign key ${FK}_{R_{2}}$ in relation $R_{1}$ then refers to the primary key PK{PK} of another relation $R_{2}$ as $\forall t\in R_{1}:~t[FK]\in\{t^{\prime}[PK]\mid t^{\prime}\in R_{2}\}\,.$ This constitutes the inter-relations in the database, with the RDBMS handling the referential integrity of ${T_{R_{1}}}[FK]\subseteq{T_{R_{2}}}[PK]$.
@@ 2.2 Graph Neural Networks
Graph Neural Networks constitute a comprehensive class of neural models designed to process graph-structured data through the concept of (differentiable) message-passing [wu2020comprehensive]. Given an input graph $G=(\mathcal{V},\mathcal{E})$, with a set of nodes $\mathcal{V}$ and edges $\mathcal{E}$, let $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}}$ be the vector representation (embedding) of node $v$ at layer $l$. The general concept of GNNs can then be defined through the following sequence of three functions:
@@ 2.3 Relational Deep Learning
In this paper, we adopt the concept of RDL as extending mainstream deep learning models, particularly the GNNs (Sec. 2.2), for application to RDBs (Sec. 2.1). For completeness, in the relational learning community [cropper2020turning30], a number of similar approaches combining relational (logic-based) and deep learning methods arose under a similar name of “deep relational learning” [vsir2021deep]. Nevertheless, for compatibility with the recently introduced frameworks [fey2024position], we hereby continue with the contemporary RDL view, where RDBs are first transformed into a graph-based representation suitable for the GNN-based learning.
@@ 2.3.1 Database Representation
The fundamental characteristic of RDL [fey2024position] is to represent an RDB as a heterogeneous graph.222sometimes referred to as the “relational entity graph” The graph representation can be defined as $G=(\mathcal{V},\mathcal{E},\mathcal{T}^{v},\mathcal{T}^{e})$, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}$ is the set of edges, $\mathcal{T}^{v}$ is a set of node types with a mapping $\phi:\mathcal{V}\to\mathcal{T}^{v}$, and $\mathcal{T}^{e}$ is a set of edge types with a mapping $\psi:\mathcal{E}\to\mathcal{T}^{e}$. The node types and edge types collectively form the graph schema $(\mathcal{T}^{v},\mathcal{T}^{e})$.
Given an RDB schema $\mathcal{R}$, the node types $T\in\mathcal{T}^{v}$ correspond to the relations (tables) $T$ within the database $\mathcal{T}^{v}\overset{1:1}{\to}\mathcal{R}$, while the edge types $\mathcal{T}^{e}$ represent the undirected inter-relations between the tables, as defined by the primary-foreign key pairs: $\mathcal{T}^{e}=\{({R_{i},R_{j}})~|~{R_{i}}[FK_{R_{j}}]\subseteq{R_{j}}[PK]~\lor~{R_{j}}[FK_{R_{i}}]\subseteq{R_{i}}[PK]\}\text{.}$ For a specific instance of an RDB $\mathcal{R}$, the set of nodes $\mathcal{V}$ is then defined as the union of all tuples (rows) $t_{i}$ from each relation $\mathcal{V}=\{v_{i,j}~|~R_{i}\in\mathcal{R},~t_{j}\in R_{i}\}\text{,}$ and the set of edges $\mathcal{E}$ is defined as $\mathcal{E}=\{({v_{i,k},v_{j,l}})|~t_{k}\in R_{i},~t_{l}\in R_{j},(R_{i},R_{j})\in\mathcal{T}^{e}\}$.
The graph representation is further enriched by node embedding matrices, attribute schema, and optionally a time mapping. Node embedding matrix $h^{(l)}_{v}\in\mathbb{R}^{d\times d_{\phi(v)}}$ contains the embedding representation of a node $v\in\mathcal{V}$ in a given layer $l$. With an attribute schema $\mathcal{A}_{T}$ that provides information about the types of attributes $A_{1},\dots,A_{n}$ associated with the nodes $v$ of a specific node type $T\in\mathcal{T}^{v}$, the initial embedding tensors $h_{v}^{(0)}\in\mathbb{R}^{d^{(0)}\times n}$ are computed from the raw database attribute tuples ${t_{i}}=(a_{1},a_{2},\ldots,a_{n})$ through multi-modal attribute encoders [fey2024position]. Finally, the time mapping is a function $\tau$ that assigns a timestamp $t_{v}$ to each node $\tau:v\mapsto t_{v}$, effectively creating a dynamically growing graph in time, enabling the use of temporal graph sampling [rossi2020temporal].
@@ 2.3.2 Predictive Tasks
In RDL, predictive tasks are implemented through the creation of dedicated training tables $T_{t}$ that extend the existing relational schema of $\mathcal{R}$. As introduced in [fey2024position], a training table $T_{t}$ contains two essential components: foreign keys $T_{t}[FK]$ that identify the entities of interest and target labels $y\in\mathcal{A}_{T_{t}}\setminus T_{t}[FK]$. Additionally, timestamps $t_{v}\in\mathcal{A}_{T_{t}}$ that define temporal boundaries for the prediction of $y$ can also be included.
@@ 2.3.3 Neural Architecture Space
    Table-level tabular model allows to employ existing tabular learning models [chen2023trompt, hu2020tabtransformer] to yield more sophisticated node embeddings $h_{v}^{(l)}$. Notably, in this stage, an RDL model may reduce the dimensionality of the node attribute matrix embedding $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}\times n}$ to a vector embedding $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}_{\phi(v)}}$.
@@
    Graph neural model then depends on the chosen embedding dimensionality of $h_{v}^{(l)}$. If there is a single embedding vector $h_{v}^{(l)}\in\mathbb{R}^{d^{(l)}_{\phi(v)}}$ per each node, the model can employ standard GNN (Sec. 2.2) heterogeneous message-passing [velivckovic2018graph, brody2022how], otherwise a custom message-passing scheme [peleska_transformers_2024] is required.
@@ 3.1 Workflow Components
The ReDeLEx workflow, depicted in Fig. 1, consists of modular blocks that enable systematic exploration of the neural architecture and database configuration space, significantly extending the current scope [robinson2024relbench] of RDL experimentation.
@@ 3.1.1 Database Connectivity
In contrast to [robinson2024relbench], the framework provides a standardized interface for connecting directly to an RDB [sqlalchemy], supporting various dialects of RDBMS. Notably, this enables a truly end-to-end deep learning pipeline, connecting to a possibly remote RDBMS, hosting the target RDB.
@@ 3.1.3 Predictive Tasks
The existing benchmark [robinson2024relbench] provides support solely for tasks with a training table $T_{t}$ generated from historical data through an SQL query. While useful, without any changes to the underlying database, this setting renders many RDB prediction tasks infeasible. ReDeLEx addresses this problem by adding support for tasks that require more substantial modifications of the original database. Tasks leveraging this functionality then not only generate a new table $T_{t}$ but a whole modified instance $\mathcal{R}^{\prime}$ of the original database $\mathcal{R}$.
@@
Importantly, this scheme can be applied to generate tasks for unsupervised pretraining. Pretraining tasks can be created by choosing any table $T\in\mathcal{R}$ in the database and duplicating it as $T^{\prime}$. The unchanged duplicate $T^{\prime}$ can then be used as a training table $T_{t}$, while the values of cells in the original table $T$ are randomly removed (masked out). The task is then to reconstruct any missing values in the classical tabular learning fashion [arik2021tabnet], opening possibilities for sophisticated pretraining methods [somepalli2021saint].
@@ 3.2 RDL-Suitable Databases and Tasks
Due to the generality of the relational model, RDBs often contain data with vastly diverse structural characteristics that, in some cases, do not properly exploit the relational model (Sec. 2.1). Likewise, not all of the 70+ available RDBs [motl2015ctu] are actually suitable for the relational learning models. In this section, we examine RDB characteristics in the context of RDL to identify suitable databases to be used in the experiments (Sec. 4).
@@ 3.2.3 Graph Data
On the other hand, RDBs are also characterized by building on the tabular representation, where an arbitrary number of attributes can be connected by a single relation. This is in contrast to the graph data which correspond to binary relational structures. Consequently, natively graph-structured data, such as molecules or family trees, although possible to be stored in an RDB, also do not fully exploit the relational model. In such cases, the RDL paradigm reduces to the simpler GNN setting [hamilton_graph_2020], introducing an unnecessary complexity otherwise. More generally, RDL models for tasks on RDBs with a low number of non-key attributes (see Sec. 4.3) may suffer from information sparsity (see App. Table 5 for databases with the stated characteristics.)